# Estimation and Analysis of Slice Propagation Uncertainty in 3D Anatomy Segmentation

### v1

@@ Abstract
Supervised methods for 3D anatomy segmentation demonstrate superior performance but are often limited by the availability of annotated data. This limitation has led to a growing interest in self-supervised approaches in tandem with the abundance of available un-annotated data. Slice propagation has emerged as an self-supervised approach that leverages slice registration as a self-supervised task to achieve full anatomy segmentation with minimal supervision. This approach significantly reduces the need for domain expertise, time, and the cost associated with building fully annotated datasets required for training segmentation networks. However, this shift toward reduced supervision via deterministic networks raises concerns about the trustworthiness and reliability of predictions, especially when compared with more accurate supervised approaches. To address this concern, we propose the integration of calibrated uncertainty quantification (UQ) into slice propagation methods, providing insights into the model’s predictive reliability and confidence levels. Incorporating uncertainty measures enhances user confidence in self-supervised approaches, thereby improving their practical applicability. We conducted experiments on three datasets for 3D abdominal segmentation using five UQ methods. The results illustrate that incorporating UQ improves not only model trustworthiness, but also segmentation accuracy. Furthermore, our analysis reveals various failure modes of slice propagation methods that might not be immediately apparent to end-users. This study opens up new research avenues to improve the accuracy and trustworthiness of slice propagation methods. Github: SlicePropUQ
@@
3D Segmentation Slice Propagation Uncertainty Quantification
@@ 1 Introduction
Achieving precise segmentation of 3D anatomy in MRI and CT volumes is critical for downstream tasks, including disease monitoring [3], diagnostic processes [12], and treatment planning [17]. Supervised deep learning models are the state of the art for 3D segmentation of various anatomies and tumors [21]. However, achieving high levels of performance requires an enormous quantity of annotated image volumes [19]. Annotating these 3D volumes, which falls outside the scope of standard clinical routines, is both costly and time-intensive, as it necessitates the specialized knowledge of several radiologists, whose expertise is often in short supply. Therefore, it is crucial to develop techniques for 3D anatomy segmentation that can operate effectively with limited or no annotations.
@@
Several machine learning approaches have been developed for 3D anatomy segmentation with limited annotations. Semi-supervised methods (e.g., [6, 25]) use a combination of fully annotated volumes as well as un-annotated volumes to train deep learning models. These methods often utilize student-teacher frameworks, enhancing model performance through mechanisms such as consistency loss or mutual information maximization. Techniques such as pseudo-labeling and sample filtering are also used to augment the annotated data pool to enhance the trained model [4, 27]. However, these methods still require high-quality, fully or partially annotated volumes for effective training and validation, limiting their utility in situations where annotated data are scarce or entirely unavailable.
@@
Slice propagation methods (e.g., [5, 30]) have evolved to offer a self-supervised approach for anatomy segmentation, eliminating the need for any annotated volumes for model training. Such methods use slice registration as a self-supervised task to establish correspondences between adjacent slices. During inference, only one annotated slice in the given volume is required to obtain a segmentation of the entire 3D anatomy. These methods reduce the burden on specialist of annotating entire volumes for training image segmentation networks. With significant amount of available un-annotated dataset, training networks on self-supervised slice-to-slice registration also significantly increases the amount of training data, enhancing the model’s ability to generalize. Moreover, these networks are trained to recognize pixel-wise correspondences of key low-level geometric features across various anatomies. As a result, they can adapt to a diverse range of anatomies, making them effective for inference on previously unseen anatomical structures. However, the performance of these self-supervised models is not yet comparable to that of supervised or semi-supervised methods trained on comparable quantities of data. Such a performance gap raises concerns about the accuracy, quality, and trustworthiness of model predictions. Given that deep learning models always produce outputs, irrespective of the confidence level of the prediction, it becomes crucial to quantify and analyze the reliability and accuracy of these self-supervised methods. This scrutiny is particularly important in clinical scenarios where false negatives and false positives can critically impact patient outcomes.
@@
Uncertainty quantification (UQ) is a key technique that helps identify when model predictions are reliable for clinical use or when extra caution is needed [14]. High predicted UQ values could flag potential incorrect segmented regions, guiding user interactions and refinement [22]. UQ analysis can also help in identifying failure modes of different methodologies that are not immediately apparent to the end-user. Uncertainty in deep learning predictions can be attributed to (a) the inherent uncertainty in the input data (aleatoric uncertainty) and (b) the uncertainty in model parameters due to limited training data (epistemic uncertainty) [16]. Aleatoric uncertainty can be directly modeled as a function of the input data. Epistemic uncertainty is more challenging to quantify, as it requires learning a distribution over model weights, but it sheds light on the model’s knowledge gaps and directly correlates with the availability and diversity of the training data. Hence, this work primarily focuses on epistemic UQ. This paper makes the following contributions to the broader goal of establishing safe and trustworthy deep-learning models for medical applications:
@@
  • A comprehensive benchmark of five state of the art methods for epistemic UQ on three datasets.
@@ 2.1 Slice Propagation Methods
Sli2Vol [30] and Vol2Flow [5] are the state-of-the-art approaches for 3D volume segmentation using a single slice annotation. Let $\mathbf{I}\in\mathbb{R}^{H\times W\times D}$ denote a 3D image, where $H$, $W$ and $D$ are the height, width, and depth of the volume. These methods provide a segmentation of the entire volume propagating one manually annotated slice $\mathbf{S}_{i}\in\mathbb{R}^{H\times W}$, where $i\in\{1,\ldots,D\}$ to the entire 3D volume by learning robust pixel-wise correspondences between adjacent slices in the volume via adjacency matrices. For example, given two adjacent slices, $\mathbf{S}_{i}$ and $\mathbf{S}_{i+1}$, the models predict an affinity matrix $\mathbf{A}_{i+1,i}$. This matrix is then used to transform $\mathbf{S}_{i}$ to acquire an estimate of slice $(i+1)$, denoted $\hat{\mathbf{S}}_{i+1}$. The difference between the estimated slice, $\hat{\mathbf{S}}_{i+1}$, and the original slice, $\mathbf{S}_{i+1}$, is then used for model training. Subsequent sections detail the differences between the methods.
@@ 2.2 Epistemic Uncertainty Quantification
Quantification of epistemic uncertainty in deep learning models is challenging since it entails learning a distribution over model weights [16]. The variance in predictions made with weights sampled from such a distribution is directly proportional to the degree of model uncertainty, where low prediction variance signifies low uncertainty or high model confidence. Although scalable epistemic UQ techniques have been proposed and successfully applied to supervised segmentation models [2], extending these advancements to slice propagation techniques has not been studied. The following section describes five state-of-the-art epistemic UQ methods and how we adapt them specifically for slice propagation tasks.
@@ 3 Results
Implementation Details. We used the original implementations of Sli2Vol [30] and Vol2Flow [5]. For UQ methods on 3D volume segmentation, we referenced the implementation release by a recent UQ benchmark [2]. We scale the slice featuring the anatomy’s largest manually annotated area as the annotated slice. We used four ensemble members for ensemble-based models and 30 posterior samples to get an average prediction for dropout and SWAG methods.  [^1] We will release our code files on GitHub at the end of the study for public use.
@@ 3.1 Uncertainty Quantification Results
Baseline deterministic models, without the integration of UQ, establish a fundamental level of accuracy, yet exhibit notable boundary detection challenges, as indicated by higher AHD values. Concrete dropout demonstrates superior segmentation and uncertainty estimation, outperforming other UQ methods in accuracy. SWAG shows higher UQ calibration than ensemble methods even with low accuracy models indicating that SWAG is a good UQ estimator, which will be important for annotation-efficient models. The SWAG method results in reduced segmentation accuracy, likely because it utilizes an average of converged weights, whereas the baseline employs the best of converged weights as assessed by the validation performance. However, SWAG provides better calibrated UQ than other methods which suggests that it is learning a wide posterior distribution of weights, resulting in diverse predictions that provide a poorer averaged estimate but accurately capture the degree of model confidence.
@@
Performance Trends. In assessing the Sli2Vol and Vol2Flow models, we also report metrics on a more granular slice-level, shedding light on critical knowledge gaps that are not captured by the mean performance statistics over dataset. Mean DSC and associated uncertainties, when averaged over large volumes, could mask the true performance nuances of these models. Our targeted analysis revealed a discernible performance drop and increased uncertainty in models as segmentation predictions deviate from the manually annotated slice, as shown in Figure 1A. This trend is reflective of a model’s diminishing accuracy and confidence as it ventures further from the annotated slice. We observe that the error and uncertainty estimates correctly correlate across distances from the annotated slice, suggesting accurate confidence estimation. Figures 1B and C bring to attention the steeper decline in surface dice scores as compared to overall DSC, signaling a quicker degradation in the model’s ability to retain surface information shortly beyond the annotated slice. This quick degradation is problematic as it indicates a rapid loss of accuracy in capturing the intricate contours and edges of anatomical structures. In medical applications, such as surgical planning or tumor resection, the precise delineation of surfaces is crucial. Imprecise surface segmentation can lead to inaccurate identification of tumor margins or critical anatomical landmarks, compromising treatment outcomes. A similar pattern emerged in the Vol2Flow analysis, shown in Figure 2, affirming this as a broader trend across both modeling approaches. Supplementary GIFs are provided to visually demonstrate the progression of predicted segmentations and associated uncertainties through a volume.
@@
Figure 1: Sli2Vol accuracy and uncertainty variation as function of distance from annotated slice A. Comparative analysis of variability in DSC and uncertainty metrics relative to the distance from the annotated slice when using concrete dropout (dataset: DecathSpleen). Performance metrics (B) DSC,(C) surface dice and (D) uncertainty for all UQ Methods, relative to the distance from the annotated slice.
Figure 2: Vol2Flow accuracy and uncertainty variation as function of distance from annotated slice A. Comparative analysis of variability in DSC and uncertainty metrics relative to the distance from the annotated slice when using MC dropout (dataset: CHAOS). Performance metrics (B) DSC,(C) surface dice and (D) uncertainty for all UQ Methods, relative to the distance from the annotated slice.
@@
Failure Modes in Slice Propagation Methods. Through detailed analysis of Figures 1 and 2, along with observations from our study, we identify several key limitations of Sli2Vol and Vol2Flow models: (a) A pronounced decline in performance metrics is observed as soon as models predict slices merely 5-20 mm away from the annotated slice. Notably, the accuracy of surface metrics diminishes starting from slices closely adjacent to the annotated slice. (b) The models exhibit difficulties in handling non-convex anatomical structures where the segmentation is discontinuous or due to branching anatomical structures. For example, across slices, femur bone segmentation becomes split into two substructures, the greater trochanter, and the femoral head. (c) Training via a surrogate registration task can inadvertently bias models toward assuming structural continuity since it emphasizes aligning continuous structures across slices. This bias results in the failure of the models to recognize the natural discontinuity or endpoints of anatomical features, resulting in over-extended segmentations.
@@ 4 Conclusion and Future Work
This study focuses on using slice propagation to reduce manual annotations in training and limit inference annotations to a single slice, addressing the trustworthiness of these methods amid minimal expert-driven supervision. We integrated five UQ methods to evaluate their accuracy and uncertainty calibration in medical image analysis. Our assessments demonstrate that incorporating UQ into slice propagation approaches enhances predictive accuracy and provides usable confidence estimation, effectively bridging the gap between semi-automatic methods and user reliance. Furthermore, our investigation uncovers critical failure modes in slice propagation methods that may go unnoticed by users. Recognizing these shortcomings is pivotal to inform improvement and continuous model refinement. Future work could explore the influence of domain variation (e.g., CT vs MRI) on uncertainty estimation. Additionally, there is substantial potential to adapt UQ techniques to more adeptly handle the intricate challenges of slice propagation. Calibrated UQ could guide the development of methodologies capable of mitigating the failure modes observed in current segmentation methods. This work reveals the potential and shortcomings of slice propagation segmentation models with UQ, increasing the potential for safe, feasible self-supervised anatomy segmentation.
@@ Acknowledgements
We thank the Scientific Computing and Imaging Institute, and the Kahlert School of Computing at the University of Utah for their support of this Project.

### v2

@@ Abstract
Supervised methods for 3D anatomy segmentation demonstrate superior performance but are often limited by the availability of annotated data. This limitation has led to a growing interest in self-supervised approaches in tandem with the abundance of available unannotated data. Slice propagation has emerged as a self-supervised approach that leverages slice registration as a self-supervised task to achieve full anatomy segmentation with minimal supervision. This approach significantly reduces the need for domain expertise, time, and the cost associated with building fully annotated datasets required for training segmentation networks. However, this shift toward reduced supervision via deterministic networks raises concerns about the trustworthiness and reliability of predictions, especially when compared with more accurate supervised approaches. To address this concern, we propose integrating calibrated uncertainty quantification (UQ) into slice propagation methods, which would provide insights into the model’s predictive reliability and confidence levels. Incorporating uncertainty measures enhances user confidence in self-supervised approaches, thereby improving their practical applicability. We conducted experiments on three datasets for 3D abdominal segmentation using five UQ methods. The results illustrate that incorporating UQ improves not only model trustworthiness but also segmentation accuracy. Furthermore, our analysis reveals various failure modes of slice propagation methods that might not be immediately apparent to end-users. This study opens up new research avenues to improve the accuracy and trustworthiness of slice propagation methods.
@@
Slice Propagation 3D Volume Segmentation Uncertainty Quantification Benchmarking Epistemic Uncertainty
@@ 1 Introduction
Achieving precise segmentation of 3D anatomy in MRI and CT volumes is critical for downstream tasks, including disease monitoring [3], diagnostic processes [12], and treatment planning [17]. Supervised deep learning models are state of the art for 3D segmentation of various anatomies and tumors [21]. However, achieving high performance levels requires an enormous quantity of annotated image volumes [19]. Annotating these 3D volumes, which falls outside the scope of standard clinical routines, is both costly and time-intensive, as it necessitates the specialized knowledge of several radiologists, whose expertise is often in short supply. Therefore, it is crucial to develop techniques for 3D anatomy segmentation that can operate effectively with limited or no annotations.
@@
Several machine-learning approaches have been developed for 3D anatomy segmentation with limited annotations. Semi-supervised methods (e.g., [6, 25]) use a combination of fully annotated volumes as well as un-annotated volumes to train deep learning models. These methods often utilize student-teacher frameworks, enhancing model performance through mechanisms such as consistency loss or mutual information maximization. Techniques such as pseudo-labeling and sample filtering are also used to augment the annotated data pool to enhance the trained model [4, 27]. However, these methods still require high-quality, fully or partially annotated volumes for effective training and validation, limiting their utility in situations where annotated data are scarce or entirely unavailable.
@@
Slice propagation methods (e.g., [5, 30]) have evolved to offer a self-supervised approach for anatomy segmentation, eliminating the need for any annotated volumes for model training. Such methods use slice registration as a self-supervised task to establish correspondences between adjacent slices. During inference, only one annotated slice in the given volume is required to obtain a segmentation of the entire 3D anatomy. These methods reduce the burden on specialists of annotating entire volumes for training image segmentation networks. With a significant amount of available unannotated datasets, training networks on self-supervised slice-to-slice registration also significantly increases the training data, enhancing the model’s ability to generalize. Moreover, these networks are trained to recognize pixel-wise correspondences of key low-level geometric features across various anatomies. As a result, they can adapt to a diverse range of anatomies, making them effective for inference on previously unseen anatomical structures. However, the performance of these self-supervised models is not yet comparable to that of supervised or semi-supervised methods trained on comparable quantities of data. Such a performance gap raises concerns about the accuracy, quality, and trustworthiness of model predictions. Given that deep learning models always produce outputs, irrespective of the confidence level of the prediction, it becomes crucial to quantify and analyze the reliability and accuracy of these self-supervised methods. This scrutiny is particularly important in clinical scenarios where false negatives and false positives can critically impact patient outcomes.
@@
Uncertainty quantification (UQ) is a key technique that helps identify when model predictions are reliable for clinical use or when extra caution is needed [14]. High predicted UQ values could flag potential incorrect segmented regions, guiding user interactions and refinement [22]. UQ analysis can also help identify failure modes of different methodologies that are not immediately apparent to the end-user. Uncertainty in deep learning predictions can be attributed to (a) the inherent uncertainty in the input data (aleatoric uncertainty) and (b) the uncertainty in model parameters due to limited training data (epistemic uncertainty) [16]. Aleatoric uncertainty can be directly modeled as a function of the input data by producing probabilistic outputs. Epistemic uncertainty, on the other hand, is more difficult to quantify as it involves learning a distribution over model weights. However, quantifying epistemic uncertainty is crucial as it highlights the model’s knowledge gaps and is directly correlated with the availability and diversity of the training data. Given that this paper focuses on weak/sparse supervision (through single-slice annotation), it is more relevant to quantify the limitations of the model’s knowledge base using epistemic uncertainty quantification (UQ). This is clinically significant for conveying the trustworthiness of the results under limited supervision. This paper makes the following contributions to the broader goal of establishing safe and trustworthy deep-learning models for medical applications:
@@
  • A comprehensive benchmark of five state-of-the-art methods for epistemic UQ on three datasets. Github Link: SlicePropUQ Github Repo.
@@ 2.1 Slice Propagation Methods
Sli2Vol [30] and Vol2Flow [5] are the state-of-the-art approaches for 3D volume segmentation using a single slice annotation. Let $\mathbf{I}\in\mathbb{R}^{H\times W\times D}$ denote a 3D image, where $H$, $W$ and $D$ are the height, width, and depth of the volume. These methods provide a segmentation of the entire volume propagating one manually annotated slice $\mathbf{S}_{i}\in\mathbb{R}^{H\times W}$, where $i\in\{1,\ldots,D\}$ to the entire 3D volume by learning robust pixel-wise correspondences between adjacent slices in the volume via adjacency matrices. For example, given two adjacent slices, $\mathbf{S}_{i}$ and $\mathbf{S}_{i+1}$, the models predict an affinity matrix $\mathbf{A}_{i+1,i}$. This matrix is then used to transform $\mathbf{S}_{i}$ to acquire an estimate of slice $(i+1)$, denoted $\hat{\mathbf{S}}_{i+1}$. The difference between the estimated slice, $\hat{\mathbf{S}}_{i+1}$, and the original slice, $\mathbf{S}_{i+1}$, is then used for model training. The following sections outline the method differences.
@@ 2.2 Epistemic Uncertainty Quantification
Quantification of epistemic uncertainty in deep learning models is challenging since it entails learning a distribution over model weights [16]. The variance in predictions made with weights sampled from such a distribution is directly proportional to the degree of model uncertainty, where low prediction variance signifies low uncertainty or high model confidence. Although scalable epistemic UQ techniques have been proposed and successfully applied to supervised segmentation models [2], extending these advancements to slice propagation techniques has not been studied. The next section describes five state-of-the-art epistemic UQ methods and their adaptation for slice propagation tasks.
@@
The selection of 5 diverse SOTA scalable epistemic UQ methods is informed by their varied approaches (covering frequentist and Bayesian perspectives) to addressing model uncertainty effectively [2]. Each method was chosen for its unique strengths: Deep and Batch Ensembles provide robustness through model averaging; MC Dropout and Concrete Dropout facilitate practical uncertainty estimation during training and inference, critical for deployment in clinical environments; and SWAG captures variability in model parameters through its approximation of the posterior distribution. This diverse toolkit allows us to comprehensively evaluate and enhance the predictive reliability and interpretability of self-supervised slice propagation methods.
@@ 3 Results
Implementation Details. We used the original implementations of Sli2Vol [30] and Vol2Flow [5]. For UQ methods on 3D volume segmentation, we referenced the implementation release by a recent UQ benchmark [2]. We scale the slice featuring the anatomy’s largest manually annotated area as the annotated slice. We used four ensemble members for ensemble-based models and 30 posterior samples to get an average prediction for dropout and SWAG methods. We use only four ensemble members based on empirical findings, balancing computational (limited GPU memory) requirements and performance gain.
@@ 3.1 Uncertainty Quantification Results
Baseline deterministic models, without the integration of UQ, establish a fundamental level of accuracy, yet exhibit notable boundary detection challenges, as indicated by higher AHD values. Concrete dropout demonstrates superior segmentation and uncertainty estimation, outperforming other UQ methods in accuracy. SWAG shows higher UQ calibration than ensemble methods, even with low-accuracy models, indicating that SWAG is a good UQ estimator, which will be important for annotation-efficient models. The SWAG method results in reduced segmentation accuracy, likely because it utilizes an average of converged weights, whereas the baseline employs the best of converged weights as assessed by the validation performance. However, SWAG provides better calibrated UQ than other methods, which suggests that it learns a wide posterior distribution of weights, resulting in diverse predictions that provide a poorer averaged estimate but accurately capture the degree of model confidence.
@@
Performance Trends. In assessing the Sli2Vol and Vol2Flow models, we also report metrics on a more granular slice level, shedding light on critical knowledge gaps that are not captured by the mean performance statistics over the dataset. Mean DSC and associated uncertainties, when averaged over large volumes, could mask the true performance nuances of these models. Our targeted analysis revealed a discernible performance drop and increased uncertainty in models as segmentation predictions deviate from the manually annotated slice, as shown in Figure 1A. This trend reflects a model’s diminishing accuracy and confidence as it ventures further from the annotated slice. We observe that the error and uncertainty estimates correctly correlate across distances from the annotated slice, suggesting accurate confidence estimation. Figures 1B and C bring to attention the steeper decline in surface dice scores as compared to overall DSC, signaling a quicker degradation in the model’s ability to retain surface information shortly beyond the annotated slice. This quick degradation is problematic as it indicates a rapid loss of accuracy in capturing the intricate contours and edges of anatomical structures. In medical applications, such as surgical planning or tumor resection, the precise delineation of surfaces is crucial. Imprecise surface segmentation can lead to inaccurate identification of tumor margins or critical anatomical landmarks, compromising treatment outcomes. A similar pattern emerged in the Vol2Flow analysis, shown in Figure 2, affirming this as a broader trend across both modeling approaches. Supplementary GIFs are provided to visually demonstrate the progression of predicted segmentations and associated uncertainties through a volume.
@@
Figure 1: Sli2Vol accuracy and uncertainty variation as a function of distance from annotated slice A. Comparative analysis of variability in DSC and uncertainty metrics relative to the distance from the annotated slice when using concrete dropout (dataset: DecathSpleen). Performance metrics (B) DSC,(C) surface dice and (D) uncertainty for all UQ Methods, relative to the distance from the annotated slice.
@@
Failure Modes in Slice Propagation Methods. Through detailed analysis of Figures 1 and 2, along with observations from our study, we identify several key limitations of Sli2Vol and Vol2Flow models: (a) A pronounced decline in performance metrics is observed as soon as models predict slices merely 5-20 mm away from the annotated slice. Notably, the accuracy of surface metrics diminishes starting from slices closely adjacent to the annotated slice. (b) The models exhibit difficulties in handling non-convex anatomical structures where the segmentation is discontinuous or due to branching anatomical structures. For example, across slices, femur bone segmentation becomes split into two substructures, the greater trochanter and the femoral head. (c) Training via a surrogate registration task can inadvertently bias models toward assuming structural continuity since it emphasizes aligning continuous structures across slices. This bias results in the failure of the models to recognize the natural discontinuity or endpoints of anatomical features, resulting in over-extended segmentations.
@@ 4 Conclusion and Future Work
This study focuses on using slice propagation to reduce manual annotations in training and limit inference annotations to a single slice, addressing the trustworthiness of these methods amid minimal expert-driven supervision. We integrated five UQ methods to evaluate their accuracy and uncertainty calibration in medical image analysis. While we do not propose a novel method of epistemic UQ estimation, we provide open-source, non-trivial extensions of existing methods to this new task and benchmark their performance. Our assessments demonstrate that incorporating UQ into slice propagation approaches enhances predictive accuracy and provides usable confidence estimation, effectively bridging the gap between semi-automatic methods and user reliance, previously unexplored. Furthermore, our investigation uncovers critical failure modes in slice propagation methods that may go unnoticed by users. Recognizing these shortcomings is pivotal to inform improvement and continuous model refinement. Future work could explore the influence of domain variation (e.g., CT vs MRI) on uncertainty estimation. Additionally, there is substantial potential to adapt UQ techniques to more adeptly handle the intricate challenges of slice propagation. Calibrated UQ could guide the development of methodologies capable of mitigating the failure modes observed in current segmentation methods. This work reveals the potential and shortcomings of slice propagation segmentation models with UQ, increasing the potential for safe, feasible self-supervised anatomy segmentation.
@@ Acknowledgements
This work was supported by the National Institutes of Health under grant numbers NIBIB-U24EB029011 and NIAMS-R01AR076120. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.